OpenStudy (anonymous):

The equation of a circle whose center is at (4, 0) and radius is length 2√(3) is (x - 4)² + y² = 2√3 (x - 4)² + y² = 12 (x + 4)² + y² = 12

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

(x-a)^2+(y-b)^2=r^2 where (a, b) center r raidus (x - 4)² + y² = 12

OpenStudy (anonymous):

thnx:)